the manager at a restaurant found that the cost to produce 150 cups of coffee is $21, while the cost to produce 300 cups of coffee is $36. assume the relationship between the cost y to produce x cups of coffee is linear.
a.) write a linear equation that expresses the cost,y, in terms of the number of cups of coffee,x.
b.) how many cups of coffee are produced if the cost of production is $56?

Answer by macston(5194)

The total cost (y) is the sum of fixed cost (F)
and the cost per cup (C) times the number of cups (x):
.
y=F+Cx
.
First, find C:
y=F+Cx
$21=F+C(150)
F=$21-C(150) Use as F
.
$36=F+C(300)
F=$36-C(300) . Replace F from above.
$21-C(150)=$36-C(300) . Add C(300) to each side.
$21+C(150)=$36 . Subtract $21 from each side.
C(150)=$15
C=$0.10 . Each cup costs $0.10 to produce.
.
Next, find fixed cost (F):
F=$21-C(150)
F=$21-$0.10(150)
F=$21-$15
F=$6 . The fixed cost is $6.
.
(a). The equation:
.
y=F+Cx.
.
y=$6+$0.10x
.
(b). For cost y=$56
.
y=$6+$0.10x
$56=$6+$0.10x
$50=$0.10x
500=x . ANSWER: 500 cups are produced for $56.